A lizard is running across the road at a speed of 2.8 m/s. To be safely on the other side of the road, the lizard needs to run 3.0 meters. How long will it take for the lizard to run 3.0 meters? Round to a tenth of a second.


1
Expert's answer
2021-02-12T04:57:42-0500

Q162865

A lizard is running across the road at a speed of 2.8 m/s. To be safely on the other side of the road, the lizard needs to run 3.0 meters. How long will it take for the lizard to run 3.0 meters? Round to a tenth of a second.


Solution:


We are given the speed with which the lizard is running and also the distance he needs to run so as to reach safely on other side of the road.


We can use the formula of speed and find the time required for the lizard to reach other side of the road.


Speed of lizard, s = 2.8 m/s

distance he needs to run = 3.0 m.


We are already given speed and distance in the SI unit, so there is no need for us to

do the conversion. Plug speed and distance in the formula.


"speed (m/s) = \frac{distance (m) }{time(s)} = \frac{d}{t}"


"2.8m/s = \frac{3.0m}{t };"


Arranging this equation for 't' we have


"t = \frac{3.0m}{2.8m/s} = 1.07 s ;"


which to tenth of a second is written as 1.1s.


Hence the lizard will take 1.1 seconds to reach safely to other side of the road.